OpenStudy (anonymous):

A box with a weight of 200N is accelerated across a level surface at 1.5 m/s^(2). What is the net force that acts on the box?

OpenStudy (shamim):

U know weight 200N is acting just downward.

OpenStudy (shamim):

We know W=mg 200=m*9.8 m=?

OpenStudy (shamim):

Now use F=ma F=m*1.5=?

OpenStudy (shamim):

Now use R^2=p^2+q^2+2pq*cos alpha R=net force=?

OpenStudy (shamim):

Given alpha=90 degree

OpenStudy (shamim):

P=200N